A person have some hens and some cows. Number of head is 38 and number of legs 100 then how many cows and hens
Found 3 solutions by stanbon, fractalier, ikleyn:
Answer by stanbon(75887)

A person have some hens and some cows. Number of head is 38 and number of legs 100 then how many cows and hens.
=======
Heads Equation:: h + c = 38
Legs Equation:: 2h +4c = 100
-------
Modify for elimination:
h + c = 38
h + 2c = 50
----
Subtract and solve for "c"::
c = 12 (# of cows)
----
h = 38-12 = 26 (# of hens)
--------------
cheers,
Stan H.
--------------

Answer by fractalier(6550)

Calls hens and cows, x and y.
Thus x + y = 38
Now when we count legs, we have
2x + 4y = 100
Let's multiply the first equation by two and subtract...
2x + 4y = 100
-(2x + 2y = 76)
which gives
2y = 24
y = 12 cows
x = 26 chickens

Answer by ikleyn(52830)

.
See the lesson  Problem on animals at a farm  in this site, where the similar problem was solved.  You will find there another method of solving.